# Low-noise Fourier Transform Spectroscopy Enabled by Superconducting On-Chip Filterbank Spectrometers

### v1

@@ Abstract
Historically employed spectroscopic architectures used for large field of view mapping spectroscopy in millimetere and sub-millimetre astronomy suffer from significant drawbacks. On-chip filterbank spectrometers are a promising technology in this respect; however, they must overcome an orders-of-magnitude increase in detector counts, efficiency loss due to dielectric properties, and stringent fabrication tolerances that currently limit scaling to resolutions of order 1000 over a large array. We propose coupling a medium-resolution Fourier transform spectrometer to a low-resolution filterbank spectrometer focal plane, which serves as a post-dispersion element. In this arrangement, medium resolution imaging spectroscopy is provided by the Fourier transform spectrometer, while the low resolution filterbank spectrometer serves to decrease the photon noise inherent in typical broadband Fourier transform spectrometer measurements by over an order of magnitude. This is achieved while maintaining the excellent imaging advantages of both architectures. We present predicted mapping speeds for a filterbank-dispersed Fourier transform spectrometer from a ground-based site and a balloon-borne platform. We also demonstrate the potential that an instrument of this type has for an line intensity mapping experiment using the James Clerk Maxwell Telescope as an example platform. We demonstrate that a filterbank-dispersed Fourier transform spectrometer would be capable of measurements of CO power spectra with a signal-to-noise ratio of 10–100 with surveys of – spectrometer hours.
@@ I Introduction
Medium resolution (.) mapping spectroscopy capable of coupling to the current and next generation large field of view (FoV) focal planes has been identified as a priority for the millimetre/sub-millimeter (mm) astronomy community [1, 2, 3, 4]. Galaxy surveys and line intensity mapping (LIM) are notable areas that stand to benefit from significant advances in instruments of this type. Specifically, the Atacama Large Aperture Submillimeter Telescope (AtLAST) is targeting a 1000 spatial pixel mapping spectrometer capable of in the 2030s to identify the “first forming galaxies” at redshifts, [5]. Currently, LIM experiments in the Mm/sub-mm experiments seek to provide competitive constraints on expansion history, neutrinos, and the dynamics of reionisation through spatially unresolved three-dimensional measurements of galaxy distributions with sensitive mapping spectrometers [6, 7, 8, 9]. An ideal LIM experiment prioritises spectrometers with high mapping speeds and large fields of view (FoVs) that target a resolving power . This has led to several mm/sub-mm LIM experiments demonstrating the capabilities of different spectrometer architectures[10, 11, 8].
@@
Historically, both grating spectrometers and Fourier transform spectrometers (FTSs) have been employed for low to medium resolution mapping spectroscopy in the mm/sub-mm. However, both spectrometer architectures suffer significant drawbacks [1]. Grating spectrometers require a spatial dimension at the focal plane to be dedicated to spectral dispersion rather than imaging and often use slit apertures that significantly reduce their throughput. FTSs have been developed for astronomical imaging spectroscopy capable of resolving powers as high as and instantaneous coverage over multiple octaves of spectral bandwidth with high throughput [12]. Additionally, FTSs have a particular multiplexing advantage (in that every detector at the focal plane measures a full spectrum) that lends itself well to imaging spectroscopy with large FoVs. In spite of these capabilities, the inherent multiplexing advantage of FTSs results in a significant increase in photon noise (being dependent on the spectral bandwidth of the light incident on a direct detector).
@@
The fabrication of FBSs also becomes a significant barrier to reaching . For mm wavelengths, an R1000 FBS would require feature variation that is less than 50 nm. While this level of accuracy has been successful in the fabrication of a single FBS, scaling these techniques across an array of several hundred to thousand FBSs has yet to be demonstrated [22].
@@ II Post-Dispersed FTS
A post-dispersed FTS combines a traditional medium- FTS with a low- dispersing spectrometer to reduce the band incident on each detector. In this arrangement, the resolution of the spectrometer is entirely determined by the FTS, while a dispersing spectrometer with an of a few hundred is capable of reducing the photon noise incident on the instrument’s detectors by over an order of magnitude (see Fig. 3). By displacing the burden of medium resolution spectroscopy to an FTS, the detector count required for such an imaging spectrometer is also reduced by an order of magnitude over that of only an FBS.
Post-dispersed FTSs employing a low-R grating spectrometer have seen some interest, leading to instrument designs and laboratory-based demonstration through the SPICA-SAFRI and PRIMA-FIRESS proposals [23, 24]. The use of an FBS as a post-dispersion spectrometer provides particular advantages over more traditional spatially dispersing architectures:
With an FBS, the spectral dispersion occurs on a transmission line that is only a few mm in length, which does not significantly scale with R. The illuminated area of a grating spectrometer scales with both R and wavelength. As such, mm/sub-mm gratings are of the order mm.
Additional focusing optics or apertures are not required, and light’s spectral content is not spatially dispersed prior to the focal plane. Therefore, there is no significant compromise to the fundamental imaging capabilities of the FTS. A grating disperser suffers from the same imaging disadvantages outlined in the previous section. Additionally, the optics required to couple an FTS to a diffraction grating result in a significant increase to the instrument’s footprint.
Traditional dispersive elements often have polarisation dependencies. As a consequence, pairs of dispersion elements may be needed to recover the full efficiency of the FTS. With an FBS disperser, polarisation can be handled at the focal plane using common polarimeter designs and techniques.
@@ III Filterbank-dispersed FTS
To demonstrate the imaging capabilities of an FBDFTS, we have calculated the mapping speed of an FBDFTS instrument from both a ground-based platform (CCAT-prime/FYST observatory) and from a balloon-borne telescope similar to the proposed BLAST observatory (see Fig. 4) [27, 28]. Other mapping spectrometers have also been included for comparison using reported values of noise equivalent flux densities, optical loads, and from their respective exposure time calculators and/or from literature (zenith pointing under ideal weather conditions) which have been averaged across their respective bands [29, 11, 30, 31]. Note that the mapping speeds for the high-resolution heterodyne instruments shown (ALMA, JCMT-HARPS) do not include the multiple retunings that would be required to cover their full spectral band (being more than 25 for each ALMA band). For the FBDFTSs, we have considered a 50% efficient FTS coupled to a 70% efficient FBS with half-wave filters. Though a single half-wave filter can be at most 50% efficient, FBSs can employ a spectral oversampling technique to improve their theoretical efficiencies to 70% [16, 20]. Also note that by employing two focal planes and separating the inputs and outputs of an FTS, it is possible to recover a near-100% efficient FTS [32]. From both platforms, an FBDFTS has the potential to provide an orders-of-magnitude improvement over a nominal FTS and an extremely competitive medium-resolution mapping spectrometer. For a ground-based FBDFTS, the CONCERTO FTS serves as a natural point of comparison, while a nominal FTS comparison has been determined for the balloon-borne FBDFTS performance.

### v2

@@ Abstract
Spectroscopic architectures that have historically been used for large field of view mapping spectroscopy in millimetere and sub-millimetre astronomy suffer from significant drawbacks. On-chip filterbank spectrometers are a promising technology in this respect; however, they must overcome an orders-of-magnitude increase in detector counts, efficiency loss due to dielectric properties, and stringent fabrication tolerances that currently limit scaling to resolutions of order 1000 over a large array. We propose coupling a medium-resolution Fourier transform spectrometer to a low-resolution filterbank spectrometer focal plane, which serves as a post-dispersion element. In this arrangement, medium resolution imaging spectroscopy is provided by the Fourier transform spectrometer, while the low resolution filterbank spectrometer serves to decrease the photon noise inherent in typical broadband Fourier transform spectrometer measurements by over an order of magnitude. This is achieved while maintaining the excellent imaging advantages of both architectures. We present predicted mapping speeds for a filterbank-dispersed Fourier transform spectrometer from a ground-based site and a balloon-borne platform. We also demonstrate the potential that an instrument of this type has for an line intensity mapping experiment using the James Clerk Maxwell Telescope as an example platform. We demonstrate that a filterbank-dispersed Fourier transform spectrometer would be capable of measurements of CO power spectra with a signal-to-noise ratio of 10–100 with surveys of – spectrometer hours.
@@ I Introduction
Medium resolution (.) mapping spectroscopy capable of coupling to the current and next generation large field of view (FoV) focal planes has been identified as a priority for the millimetre/sub-millimeter (mm) astronomy community [16, 5, 30, 45]. Galaxy surveys and line intensity mapping (LIM) are notable areas that stand to benefit from significant advances in instruments of this type. Specifically, the Atacama Large Aperture Submillimeter Telescope (AtLAST) is targeting a mapping spectrometer with 1000 spatial pixels capable of resolving powers in the 2030s to identify the “first forming galaxies” at redshifts, [31]. Currently, LIM experiments in the mm/sub-mm experiments seek to provide competitive constraints on expansion history, neutrinos, and the dynamics of reionisation through spatially unresolved three-dimensional measurements of galaxy distributions with sensitive mapping spectrometers [33, 48, 28, 29]. An ideal LIM experiment prioritises spectrometers with high mapping speeds and large fields of view (FoVs) that target a resolving power . This has led to several mm/sub-mm LIM experiments demonstrating the capabilities of different spectrometer architectures[56, 12, 28].
Historically, both grating spectrometers and Fourier transform spectrometers (FTSs) have been employed for low to medium resolution mapping spectroscopy in the mm/sub-mm. However, both spectrometer architectures suffer significant drawbacks [16].
@@
FTSs have been developed for astronomical imaging spectroscopy capable of resolving powers as high as with high throughput [20]. Additionally, FTSs provide access to a broad free spectral range (FSR), in practice being only limited by the performance of the optics employed (often the beamsplitter). Thus, in the mm/sub-mm, FTSs covering multiple octaves of spectral bandwidth are achievable [21, 4]. Also of note for wide-field imaging spectroscopy, FTSs have a particular multiplexing advantage in that every detector at the focal plane measures a full spectrum. In spite of these capabilities, the inherent multiplexing advantage of FTSs results in a significant increase in photon noise (being dependent on the spectral bandwidth of the light incident on a direct detector).
If the inherent noise of an FTS from spectral multiplexing can be reduced, it holds significant advantages over grating spectrometers, which struggle to scale to in the mm/sub-mm. The throughput disadvantage as a consequence of the slit apertures employed by gratings at medium to high resolution is thoroughly examined by Jacquinot (1960) [26]. Here we note that the throughput of a spectrometer limited only by a circular input aperture, such as an FTS, is given as
where is the effective area of the interferometer beam and is the resolving power of the spectrometer, while for an optimal slit-fed grating spectrometer it is given as
where is the effective area of the beam in the grating spectrometer, is the angular height of the slit, and and are the angles of incidence and emergence, respectively. From these, we see that the throughput of an interferometric spectrometer exceeds that of such a grating spectrometer (with equal area and resolution) by over an order of magnitude for all practical physical dimensions of such a grating. It should also be noted that spectrum from a diffraction grating is dispersed along a spatial dimension of a focal plane, making it challenging to couple a grating to a large FoV [16].
Fabry-Perot Spectrometers (FPSs), in a similar manner to FTSs, have a throughput advantage over grating spectrometers. Additionally they are easily capable of resolving powers of order . However, FPS are inherently narrow banded as increasing the FSR of an FPS with a given resolution requires increasing the finesse of the resonant cavity. In an ideal FPS, finesse is dictated by the reflectivity of the cavity. However, it should be noted that the acceptance of off-axis rays from a large FoV and defects in the cavity result in a diminished throughput that scales with the reflective finesse of the cavity [1]. In practice, FPSs have a relatively narrow instantaneous free spectral range, being of order 1/20th of an octave at and scaling with . As a consequence, FPSs require a spectral scanning mechanism to cover an appreciable spectral range and the precision of the actuation mechanism becomes a significant restriction on an FPS [16].
Here we note an important subtlety; though both FTSs and FPIs employ a scanning mechanism, in an FTS all spectral components are measured at all times. Additionally, the scanning mechanism in an FTS dictates the maximum spectral resolution that can be achieved in an FTS. FTS scanning mechanisms with sufficient performance for have substantial heritage from terrestrial, balloon, and space based platforms [18, 22, 10].
@@
The fabrication of FBSs also becomes a significant barrier to reaching . In a typical lumped-element arrangement, the central frequency of FBS spectral channels is largely dictated by the length of the half-wave filter [46]. Thus, this imposes a limit on fabrication tolerances to develop a spectrometer that can effectively and efficiently sample a given spectral bandpass. For mm wavelengths, an R1000 FBS would require feature variation that is less than 50 nm [54]. While this level of accuracy has been successful in the fabrication of a single FBS, scaling these techniques across an array of several hundred to thousand FBSs has yet to be demonstrated.
@@ II Post-Dispersed FTS
A post-dispersed FTS combines a traditional medium- FTS with a low- dispersing spectrometer to reduce the band incident on each detector. In this arrangement, the resolution of the spectrometer is entirely determined by the FTS, while a dispersing spectrometer with an of a few hundred is capable of reducing the photon noise incident on the instrument’s detectors by over an order of magnitude. This is demonstrated in Fig. 3 comparing against two representative FTS bands. Here we note that a single FBS focal plane employing lumped-element half-wave resonators is capable a full octave of bandwidth with band centres chosen to reduce load from atmospheric features [2]. The optical load used in this Fig. 3 determined using a model of a 15-m ground-based telescope coupled to an FTS with warm optics under a 25th percentile weather condition at Maunakea. A 70% efficient backend dispersing spectrometer has been chosen as it is representative of the maximum theoretical efficiency of half-wave FBS and of similar efficiency to current generation sub-mm grating spectrometers [56, 3]. In addition to this reduction in noise, displacing the burden of medium resolution spectroscopy to an FTS reduces the detector count required for such an imaging spectrometer by an order of magnitude over that of only an FBS array.
Post-dispersed FTSs employing a low-R grating spectrometer have seen some interest, leading to instrument designs and laboratory-based demonstration through the SPICA-SAFARI and PRIMA-FIRESS mission proposals/studies [38, 8, 7]. In particular, Bradford et al. (2025)[7] note the advantages a post-dispersed FTS instrument has over other spectrometer designs with similar performance targets (including FPS-based designs). The use of an FBS as a post-dispersion spectrometer provides particular advantages over more traditional spatially dispersing architectures:
With an FBS, the spectral dispersion occurs over a total footprint that is about a centimetre in length at while the illumination area of a grating spectrometer scales significantly with wavelength. As such, mm/sub-mm gratings of similar resolution are of the order 1 m in their size. Additionally, coupling an FTS to a grating requires additional optics (including an intermediate focus), significantly increasing the size and optical complexity of the post-dispersed FTS [44].
Light’s spectral content is not spatially dispersed prior to the focal plane. Therefore, there is no significant compromise to the fundamental imaging capabilities of the FTS. A grating disperser suffers from the same imaging disadvantages outlined in the previous section. Again, we note that a grating spectrometer sacrifices a spatial-imaging dimension in order to measure spectral content. In principle, an FPS can displace spectral dispersion away from a 2-dimensional imaging array of antennae (see Fig. 4 and Barry et al 2022 [2]).
Traditional dispersive elements often have polarisation dependencies. As a consequence, pairs of dispersion elements may be needed to recover the full efficiency of the FTS. With an FBS disperser, polarisation can be handled at the focal plane using common polarimeter designs and techniques. E.g., coupling a dual-polarisation orthomode transducer antenna to two filterbanks, one for each polarisation [2, 57].
@@ III Filterbank-dispersed FTS
To demonstrate the imaging capabilities of an FBDFTS, we have calculated the mapping speed of an FBDFTS instrument from both a ground-based platform (CCAT-prime/FYST observatory) and from a balloon-borne telescope similar to the proposed BLAST observatory (see Fig. 5) [11, 13]. Other mapping spectrometers have also been included for comparison using reported values of noise equivalent flux densities and optical loads from their respective exposure time calculators and/or from literature (zenith pointing under ideal weather conditions) [25, 12, 6, 11]. We note that for EoR-Spec, its instantaneous bandwidth is marked by vertical bars and the time penalty required to measure its full band has been included in the calculation. For the FBDFTSs, we have considered a 50% efficient FTS coupled to an FBS with half-wave filters. Though a single half-wave filter can be at most 50% efficient, FBSs can employ a spectral oversampling technique to improve their theoretical efficiencies to 70% [3, 46]. From both platforms, an FBDFTS has the potential to provide an orders-of-magnitude improvement over a nominal FTS and an extremely competitive medium-resolution mapping spectrometer, reaching similar mapping speeds to low-resolution designs. For a ground-based FBDFTS, the CONCERTO FTS serves as a natural point of comparison, while a nominal FTS comparison has been determined for the balloon-borne FBDFTS performance. Of particular note, an FBDFTS has the potential of providing similar mapping speeds to EoR-Spec with an order of magnitude improvement in spectral resolution. To achieve a similar level of resolution with the EoR-Spec FPS, the instantaneous bandwidth would also be reduced by the same factor.
We note some additional subtleties that can arise from separating the dual output ports of an FTS as in the Mach-Zehnder design [40]. Since such an FTS naturally lends itself well to two focal planes, dramatic improvements over the projected performance in Fig. 5 and/or practical improvements beyond what we have discussed can be obtained. (1) Employing a duplicate FBS focal plane at the second FTS output, would make it possible to recover a near-100% efficient FTS. (2) A second FBS focal plane designed around a different band could be used to increase the spectral bandwidth of the instrument to two octaves [22]. (3) The second import can be place off-axis on the sky to provide natural background subtraction [19]. (4) Polarisation could be separated in the FTS’s interferometer portion, terminating on separate focal planes dedicated to orthogonal polarisation states [37]. In many of these and similar designs, one could consider a single focal plane instrument as baseline/demonstrator that could later be upgraded.